OpenStudy (mayankdevnani):

So CONFUSING FOR ME ??? Which of the following is greatest ? A) cosec 1 B) cosec 2 C) cosec 4 D) cosec (-6)

OpenStudy (anonymous):

\[\csc (1) = \frac{ 1 }{ \sin (1) }\] \[csc (2) = \frac{ 1 }{ sin (2) }\] \[csc (4) = \frac{ 1 }{ sin (4) }\] \[csc (-6) = \frac{ 1 }{ sin (-6) }\]

OpenStudy (mayankdevnani):

correct !

OpenStudy (mayankdevnani):

so , sin 4 > sin 2 > sin 1> sin (-6) is it correct ??

OpenStudy (mayankdevnani):

@dimensionx

OpenStudy (anonymous):

yes, sin 4 > sin 2 > sin 1 > sin (-6)

OpenStudy (mayankdevnani):

so , cosec (-6) is our answer

OpenStudy (anonymous):

Well, csc (-6) will yield a negative value

OpenStudy (mayankdevnani):

yeah...so this is so small

OpenStudy (mayankdevnani):

therefore, cosec 1 is our answer

OpenStudy (anonymous):

bingo!

OpenStudy (mayankdevnani):

but this is not the answer !

OpenStudy (anonymous):

why?

OpenStudy (mayankdevnani):

lol.. i don't know that's why i am asking here ?

OpenStudy (anonymous):

it should be...

OpenStudy (mayankdevnani):

i hope you are correct !

OpenStudy (anonymous):

sin(1) is the smallest (other than sin(-6) which is negative), so 1/sin(1) should be the greatest.

OpenStudy (mayankdevnani):

we should take some suggesstions from the best users !

OpenStudy (mayankdevnani):

@ganeshie8 @e.mccormick @eliassaab @Whitemonsterbunny17 @Destinymasha @jim_thompson5910 @dan815 @Callisto

OpenStudy (mayankdevnani):

@texaschic101

OpenStudy (mayankdevnani):

@satellite73

OpenStudy (anonymous):

Well what is the correct answer?

OpenStudy (callisto):

Are the angles in degree or radian?

OpenStudy (mayankdevnani):

i don't know ! suppose they are in degree

OpenStudy (callisto):

If they are in degrees , you are right; if they are in radians, you are wrong

OpenStudy (mayankdevnani):

yeah !!

OpenStudy (mayankdevnani):

\[\large \bf \sin(1^c)=\sin(\frac{180}{\pi})\]

OpenStudy (mayankdevnani):

and this is approximate to :- 0.679

OpenStudy (mayankdevnani):

\[\large \bf \sin(2^c)=\sin(\frac{2 \times 180}{\pi})=0.997\]

OpenStudy (callisto):

A) cosec 1 = 1/(sin57) B) cosec 2 = 1/(sin 114) = 1/sin(66) C) cosec 4 = 1/(sin 229) = something -ve D) cosec (-6) = 1/(sin -343) = 1/sin(16)

OpenStudy (mayankdevnani):

D) cosec (-6) = 1/sin(-342)=1/sin (18)

OpenStudy (anonymous):

Well in that case,

OpenStudy (callisto):

sin(66) > sin(57) > sin(16) => csc(66) < csc(57) < csc(16)

OpenStudy (mayankdevnani):

yeah !! now i understand

OpenStudy (mayankdevnani):

thank you @Callisto and @dimensionx

OpenStudy (mayankdevnani):

@dimensionx OPTION D IS CORRECT ANSWER

OpenStudy (mayankdevnani):

but how do we know that value is in degree or radian ? like in this question

OpenStudy (mayankdevnani):

@Callisto

OpenStudy (mayankdevnani):

if we choose degree,then option A is correct if we choose radian,then option D is correct

OpenStudy (callisto):

According to my teacher, in advanced mathematics, especially in calculus, we don't use degree

OpenStudy (mayankdevnani):

see,if i tick option A,then my answer is wrong according to answer key because i take degree

OpenStudy (mayankdevnani):

oh !! i see @Callisto

OpenStudy (mayankdevnani):

thanks for your help !

OpenStudy (callisto):

@mayankdevnani Your claim that "D) cosec (-6) = 1/sin(-342)=1/sin (18)" is not quite true. Maybe you should check your calculation.

OpenStudy (mayankdevnani):

\[\large \bf 57 \times (-6)=-342\]

OpenStudy (mayankdevnani):

am i wrong here ??

OpenStudy (mayankdevnani):

@Callisto

OpenStudy (mayankdevnani):

are you there ?????

OpenStudy (callisto):

\(\frac{180}{\pi}\times (-6) = -343.7746771...\)

OpenStudy (mayankdevnani):

i take approximation. since,we know that 1 radian =57.295 so, 57.295(-6)=-343.77

OpenStudy (mayankdevnani):

i have another question ;- if centre of circle is \(\large \bf (\pi,e)\),then how many rational points will lie on circle ?

OpenStudy (mayankdevnani):

@Callisto and @dimensionx and @ganeshie8

OpenStudy (mayankdevnani):

last one !

OpenStudy (callisto):

Make a new post for a new question

OpenStudy (mayankdevnani):

okies

OpenStudy (anonymous):

In radians \[ \csc (-6)>\csc (1)>\csc (2)>\csc (4) \] In degrees \[ \csc (1 {}^{\circ})>\csc (2 {}^{\circ})>\csc (4 {}^{\circ})>\csc (-6 {}^{\circ}) \]